# Adaptive Interactive Navigation of Quadruped Robots using Large Language Models

### v1

@@ III-B Skill Pre-training with Reinforcement Learning
With low-level locomotion policies, we train high-level skills using the hierarchical RL method as a pose-tracking task, which generates velocity commands for the pre-trained locomotion policy to follow. Specifically, the pushing skill aims to push boxes to the target pose, and the navigation skill navigates the robot to the target pose while avoiding obstacles. We randomize the shape and size of the boxes for pushing skill learning and obstacle density for navigation skill learning, facilitating robust high-level skill learning. All these skills are trained using PPO [21] in the IsaacLab simulation environment. After training these primitive skills, we encapsulate them into APIs for task planning, enabling sequential skill execution to accomplish complex tasks.
III-C Adaptive Replanning with Observation Interpretation
@@
Unlike previous LLM-based planning approaches that assume omniscient scene description, navigation tasks typically involve unknown environments, where the robot continuously acquires new observations with an egocentric perspective. For this reason, we develop the adaptive replanning mechanism to analyze the new observations and efficiently replan based on new environmental comprehension. This mechanism consists of two LLM-based roles: an advisor and an arborist, designed to address two key questions: when to replan, and how to enhance replanning efficiency.
@@
The advisor analyzes new environmental information from visual perception and the current plan to determine whether replanning is necessary based on these comprehension updates. We let the advisor conduct an environmental interpretation in three aspects, including failure, new objects, and revaluation. Failure indicates that the current plan has encountered some errors, new objects represent the discovery of a new interactive object that may be useful, and revaluation means reassessing the plan based on updated information about the environment. Unlike previous works that only replan based on failure, the other two types of comprehension are also important to perform tasks with an incrementally updated understanding of the environment. For example, when you discover a new object, you need to analyze its geometric and semantic information and decide to utilize it if it is more helpful to your task. As for revaluation, the robot’s understanding of the object information under observation uncertainties will be updated, which will affect the evaluation of the current plan. Based on these three types of comprehension, the advisor will decide whether to replan and replan suggestions.
@@
When the advisor determines that replanning is necessary, it generates specific suggestions for the arborist, which subsequently modifies the primitive tree structure through node addition and pruning operations. Node addition incorporates new object information, while node pruning removes failed paths. This tree-based architecture enables computationally efficient replanning through its inherent structural flexibility, facilitating real-time task execution. Following structural modifications, the system performs backup operations within the updated tree and selects an optimal plan skeleton for execution using pre-trained skills. Through the adaptive replanning approach, the robot can respond to new environmental information flexibly and adjust the planning strategies accordingly. By analyzing new observations, the robot can assess their impact on task completion and decide whether adjustments to the current plan are required. This process helps avoid the computational cost of unnecessary replanning.
IV Simulation
In this section, we present the simulation experiments conducted to validate the effectiveness of the proposed method. The experiments are structured into four main parts: (1) the experimental setup, which outlines the simulation environment and interactive navigation tasks; (2) the baseline methods compared to AINav to establish a performance benchmark; (3) the simulation results and analysis; (4) ablation studies to investigate the contribution of individual components of the proposed method.
IV-A Experimental Setup
Simulation Settings. To validate the effectiveness of our proposed method, we conduct extensive simulations in IsaacLab. The simulations are conducted in a $8$m $\times$ $8$m space with fixed obstacles and movable boxes. We design diverse simulation environments corresponding to four interactive navigation tasks, which will be detailed in the following section. We use the Unitree Go2 as the robotic platform to perform the simulations. All experiments are conducted on a desktop with an Intel Core i7 CPU @ 2.10 GHz and 16GB RAM.
Task Description. We design four interactive navigation tasks: box obstruction, box usage, stair building, and an integrated task, as illustrated in Fig. 3. The box obstruction task involves environments where boxes block the path. Box usage and stair building require the robot to interact with boxes to complete navigation tasks. The integrated task combines challenges from the previous three tasks. Each scenario is divided into three levels of difficulty: low, medium, and high. The low-difficulty scenario can be regarded as the traditional navigation task where the goal point can be reached without interaction. Medium difficulty requires an interactive environment, such as pushing movable obstacles or using a box as a tool, while high difficulty involves additional challenges with immovable boxes, designed to evaluate the replanning performance and efficiency of various methods. The robot needs to reach the designated goal, with a maximum climbing capability of $0.3$ meters. We set the maximum simulation time to $120$s and conducted $10$ repeated experiments for each scenario.
Evaluation Metrics. The evaluation metrics include success rate (SR), overall time (OT), overall time under successful trials (OTS), planning time (PT), and execution time (ET) to comprehensively assess the performance of different methods. SR refers to the probability of the robot successfully reaching the target point across 10 trials. OT denotes the total simulation time across all scenarios. If a trial fails, OT is recorded as the maximum simulation time, which is $120$s. OTS represents the overall completion time for successful task completion. PT and ET indicate the time spent on task planning and skill execution under successful completion, respectively. SR, OT, and OTS indicate the comprehensive effectiveness and efficiency of methods, while planning time and execution time reflect the computational efficiency and quality of task planning.
Figure 3: Interactive tasks in simulations.
IV-B Baseline Methods
We compare the performance of the proposed method to the following baselines. Art-planner [11] is a sampling-based approach with learned motion cost for quadruped robots, chosen as a representative approach to solve traditional navigation tasks. Hierarchical RL trains a high-level policy to function as the task planning module, which decides which low-level skill to use and the corresponding parameters. The low-level skills utilized in hierarchical RL are identical to those in our method. The reward is inversely related to the distance between the robot and the goal, encouraging the robot to move toward the target. Furthermore, we also select two LLM-based method for comparison. The first is RoboTool [6], which employs multiple LLMs to perform different roles and invoke low-level skills to utilize tools, thereby solving complex planning and reasoning challenges. Since this method uses open-loop formulation without reaction to new objects, we provide the complete environment description as its input and execute the skills suggested by its output. The second method is Tree-planner [7], which employs a tree-structured reasoning approach to solve closed-loop TAMP problems. Because its plan sampling module requires global information, we include the initial object information as input. We adjust the prompts used in these two LLM-based methods for better adaptation to the interactive navigation tasks. For a fair comparison, all methods use the same set of low-level skills as our method.
IV-C Results and Analysis
Simulation results are displayed in Table I. We can observe that our method exhibits higher success rates and less overall time compared to other baselines. The Art-planner achieves high task success rates in low-difficulty scenarios with low execution time, showcasing its effective planning capabilities in traditional navigation tasks. However, in medium-difficulty and high-difficulty scenarios that require interaction, Art-planner is limited to passively adapting to the current environment without the capability to actively alter the workspace, thereby failing to accomplish more complex interactive navigation tasks. Hierarchical RL can achieve limited success in low-difficulty scenarios of box obstruction, while in more intricate tasks that require interaction to solve, hierarchical RL struggles to identify the precise sequence of skills necessary for task completion in the expansive search space.
@@ IV-C Results and Analysis
Method Metric Box obstruction Box usage Stair building Integrated task
L M H L M H L M H L M H
Art-planner SR $\uparrow$ 100% 0% 0% 80% 0% 0% 90% 0% 0% 80% 0% 0%
OT $\downarrow$ 33.3 120.0 120.0 50.2 120.0 120.0 44.0 120.0 120.0 57.2 120.0 120.0
OTS $\downarrow$ 33.3 - - 32.8 - - 35.6 - - 41.5 - -
PT $\downarrow$ 5.6 - - 5.9 - - 5.8 - - 6.7 - -
ET $\downarrow$ 27.7 - - 26.9 - - 29.8 - - 34.8 - -
@@
Robotool SR 90% 80% 30% 90% 70% 20% 80% 60% 0% 60% 30% 0%
OT 65.9 79.5 104.3 64.8 80.5 109.2 70.9 92.5 120.0 88.0 108.7 120.0
OTS 59.9 69.4 67.8 58.7 63.6 66.2 58.6 74.1 - 66.6 82.2 -
PT 25.6 27.2 24.9 28.3 26.4 28.8 26.1 29.3 - 23.5 25.7 -
ET 34.3 42.2 42.9 30.4 37.2 37.4 32.5 44.8 - 43.1 56.5 -
Tree-planner SR 90% 80% 60% 90% 50% 40% 80% 40% 30% 70% 30% 0%
OT 49.7 65.6 89.5 53.4 92.9 104.9 72.9 100.6 115.1 90.3 116.6 120.0
OTS 41.9 52.0 69.1 46.0 65.8 82.2 61.1 71.5 103.6 77.6 108.5 -
PT 10.7 15.7 18.8 21.2 25.9 29.8 26.4 28.5 36.1 29.3 36.2 -
ET 31.2 36.3 50.3 24.8 39.9 52.4 34.7 43.0 67.5 48.3 72.3 -
Ours SR 100% 100% 90% 90% 70% 60% 90% 80% 50% 80% 60% 50%
OT 42.3 53.3 77.4 50.0 78.7 96.7 52.6 85.3 108.5 69.7 98.9 113.8
OTS 42.3 53.3 72.7 42.2 61.0 81.1 45.1 76.6 96.9 57.1 84.9 107.5
PT 14.1 19.4 30.1 15.7 26.4 36.2 14.8 29.7 40.2 20.4 34.1 39.2
ET 28.2 33.9 42.6 26.5 34.6 44.9 30.3 46.9 56.7 36.7 50.8 68.3
  • L = Low difficulty, M = Medium difficulty, H = High difficulty
  • SR = Success Rate, OT = Overall time, PT = Planning Time, ET = Execution Time
@@
Figure 5: Success cases of the proposed method in four interactive navigation tasks. The red phases indicate the robot interacts with objects, while the green phases represent the robot performing replanning based on new environmental observations.
@@
Leveraging the scene understanding and reasoning capabilities of pre-trained LLM, Robotool and tree-planner achieve a higher success rate in interactive navigation tasks compared to traditional navigation methods. However, Robotool is constrained by its open-loop planning scheme, which prevents it from adapting to new environmental information, resulting in lower success rates in high-difficulty tasks. While tree-planner leverages a closed-loop design to adjust its decisions based on new observations, its performance is highly dependent on the quality of the sampled plans, resulting in instability in planning results. Specifically, once the tree structure is generated, its inflexibility can hinder the planner from obtaining the optimal solution that does not exist in the original tree or changing to an updated better strategy in response to new observations. Additionally, both methods require omniscient environmental information to generate task decomposition plans, rendering them ineffective in unknown environments. In contrast, our method not only enhances the quality of task planning outcomes through the design of a proposer and evaluator but also leverages adaptive replanning to flexibly adapt to new environmental information. This ensures efficient planning and real-time responsiveness in unfamiliar environments without global environmental information as prior, resulting in high success rates in various scenarios, especially in challenging situations that require replanning. Although node evaluation incurs additional planning time, the improved task planning quality and flexible replanning mechanism enable our method to achieve higher success rates and shorter overall completion time compared to other LLM-based approaches. As shown in Fig. 4, AINav demonstrates the highest success rate and the lowest overall completion time across various scenarios, showcasing its superior effectiveness in complex navigation tasks.
@@
We also present successful cases of our method across four tasks in Fig. 5. In the box obstruction scenario, the robot demonstrates the ability to comprehend the blocked environment, move boxes to create free space to reach the target position. In the box usage and stair building scenarios, the robot actively utilizes movable obstacles to construct stairs, enabling access to higher areas, which is unreachable for a quadruped robot under normal circumstances. In integrated task scenario, when confronted with various potential interaction tasks, the robot understands the current issue by VLM and intelligently interacts with the appropriate objects to facilitate task completion. Furthermore, we can observe that the robot is capable of responding to updated information, such as when the box fails to move or the robot identifies new objects and tasks. By leveraging an efficient closed-loop system, the proposed approach enhances the robot’s capability to adapt to incremental environmental comprehension, thereby facilitating effective task completion.
IV-D Ablation Studies

### v2

@@ IV-C Skill Pre-training with Reinforcement Learning
Building upon low-level locomotion policies, we train high-level skills using the hierarchical RL method as a pose-tracking task, which generates velocity commands for the pre-trained locomotion policy to follow. Specifically, the pushing skill aims to push boxes to the target pose, and the navigation skill navigates the robot to the target pose while avoiding obstacles. We randomize the shape and size of the boxes for pushing skill learning and obstacle density for navigation skill acquisition, facilitating robust high-level skill learning. All these skills are trained using PPO [schulman2017proximal] in the IsaacLab simulation environment. The reward terms of all skills are detailed in Section˜-B. After training these primitive skills, we encapsulate them into APIs for task planning, enabling sequential skill execution to accomplish complex tasks.
IV-D Adaptive Replanning with Observation Interpretation
@@
Unlike previous LLM-based planning approaches that assume omniscient scene description, navigation tasks typically involve unknown environments, where the robot acquires new observations with an egocentric perspective. For this reason, we develop the adaptive replanning mechanism to analyze the new observations and efficiently replan when needed. This mechanism consists of two LLM-based agents, an advisor and an arborist, designed to address two key questions: when to replan, and how to enhance replanning efficiency.
@@
Taking new environmental observations and the current plan as inputs, the advisor conducts an observation interpretation and determines whether replanning is necessary. We propose three aspects of observation interpretation, including failure, new objects, and revaluation. Failure indicates the current plan is practically infeasible, new objects represent the perception of a new object that may be useful, and revaluation means reassessing the plan based on updated environmental information. Unlike previous works that only replan when encountering a failure situation, the other two types of comprehension are also important to perform tasks with an incrementally updated understanding of the environment. For example, upon detecting a new object, the robot needs to assess its geometric properties and determine whether utilizing it would enhance task performance. As for revaluation, perceived object properties, such as whether movable, may change during the robot’s movement, leading to updated affordances that may necessitate revaluating the interaction strategies in the current plan. Triggered by one of these three interpretations, the advisor decides whether to replan by comparing potential improvements against the current plan.
@@
If the advisor determines that replanning is necessary, it generates specific suggestions about adjustments to the current plan according to the interpretation of the new observations for the arborist, which subsequently modifies the primitive tree structure through node addition and pruning operations. Node addition takes new objects into consideration in the primitive tree, while node pruning removes failed paths. This tree-based architecture enables computationally efficient replanning through its inherent structural flexibility, facilitating real-time task execution. Following structural modifications, the system performs backpropagation operations within the updated tree and selects an optimal plan skeleton for execution using pre-trained skills. Through the adaptive replanning approach, the robot can respond to new environmental information flexibly and adjust the planning strategies accordingly.
V Simulation
Figure 4: Simulation environments for interactive navigation tasks.
V-A Experimental Setup
Simulation Settings. We conduct simulations in IsaacSim to validate the effectiveness of AINav. The simulation environment includes the maze, indoor Matterport3d dataset, urban ruin, and field ruin, as shown in Fig.˜4. In the maze environment, we design diverse tasks and scenarios, which will be detailed in the following section. For the other three environments, we test AINav in simulated realistic indoor and outdoor scenarios, including cluttered domestic environments and ruin scenes. We use the GPT-4o model as the LLM backbone for all methods in our experiments, the CLIP model as VLM, the YOLOv8 model as the detection model, and the Unitree Go2 as the robotic platform to perform the simulations. To simplify visual processing, we assume that the robot can obtain noisy dimensions of objects within the robot’s field of view, modeled as white noise with a standard deviation proportional to the distance between the robot and the objects. All simulations are conducted on a desktop with an Intel Core i7 CPU @ 2.10 GHz and 16GB RAM.
@@ V-A Experimental Setup
Method Metric Box obstruction Box usage Stair building Matterport Urban ruin Field ruin
L M H L M H L M H
Art-planner SR $\uparrow$ 100% 0% 0% 80% 0% 0% 90% 0% 0% 0% 0% 0%
OT $\downarrow$ 33.3 120.0 120.0 50.2 120.0 120.0 44.0 120.0 120.0 120.0 120.0 120.0
OTS $\downarrow$ 33.3 - - 32.8 - - 35.6 - - - - -
PT $\downarrow$ 5.6 - - 5.9 - - 5.8 - - - - -
ET $\downarrow$ 27.7 - - 26.9 - - 29.8 - - - - -
TLS $\downarrow$ 14.8 - - 17.4 - - 16.8 - - - - -
@@
TLS 24.1 - - - - - - - - - - -
RoboTool SR 90% 80% 30% 90% 70% 20% 80% 60% 0% 30% 0% 20%
OT 65.9 79.5 104.3 64.8 80.5 109.2 70.9 92.5 120.0 110.3 120.0 113.9
OTS 59.9 69.4 67.8 58.7 63.6 66.2 58.6 74.1 - 87.8 - 89.7
PT 25.6 27.2 24.9 28.3 26.4 28.8 26.1 29.3 - 23.5 - 28.4
ET 34.3 42.2 42.9 30.4 37.2 37.4 32.5 44.8 - 64.3 - 61.3
TLS 19.6 24.3 29.5 19.2 23.7 26.7 20.9 31.4 - 43.4 - 36.7
Tree-planner SR 90% 80% 60% 90% 50% 40% 80% 40% 30% 50% 0% 0%
OT 49.7 65.6 89.5 53.4 92.9 104.9 72.9 100.6 115.1 103.6 120.0 120.0
OTS 41.9 52.0 69.1 46.0 65.8 82.2 61.1 71.5 103.6 87.2 - -
PT 10.7 15.7 18.8 21.2 25.9 29.8 26.4 28.5 36.1 32.2 - -
ET 31.2 36.3 50.3 24.8 39.9 52.4 34.7 43.0 67.5 55.0 - -
TLS 19.1 22.8 34.6 16.7 28.5 34.0 22.3 24.9 46.1 38.4 - -
AINav (Ours) SR 100% 100% 90% 90% 70% 60% 90% 80% 50% 90% 30% 50%
OT 42.3 53.3 77.4 50.0 78.7 96.7 52.6 85.3 108.5 77.3 113.1 97.8
OTS 42.3 53.3 72.7 42.2 61.0 81.1 45.1 76.6 96.9 72.7 97.1 75.6
PT 14.1 19.4 30.1 15.7 26.4 36.2 14.8 29.7 40.2 19.5 28.2 30.8
ET 28.2 33.9 42.6 26.5 34.6 44.9 30.3 46.9 56.7 53.2 68.9 44.8
TLS 17.4 18.9 25.6 16.2 19.2 26.4 17.1 23.4 31.3 34.9 24.6 21.5
  • L = Low difficulty, M = Medium difficulty, H = High difficulty, SR = Success rate, OT = Overall time, OTS = Overall time under successful trials, PT = Planning time, ET = Execution time, TLS = Trajectory length under successful trials
Task Description. In the maze environment, we design three interactive navigation tasks: box obstruction, box usage, and stair building, as illustrated in Fig.˜6. The box obstruction task involves environments where boxes block the path. Box usage and stair building require the robot to utilize boxes as tools to complete navigation tasks. Each scenario is divided into three levels of difficulty: low, medium, and high. The low-difficulty scenario removes the interaction tasks for simple navigation, which can be regarded as traditional navigation. Medium difficulty requires an interaction, such as pushing movable obstacles or using a box as a tool, while high difficulty involves additional challenges with heavy boxes that are difficult to push, designed to evaluate the replanning performance and efficiency of various methods. In the Matterport3d and ruin environments, we designate the start and navigation goal shown in Fig.˜7 to test the effectiveness of AINav in diverse environments. The robot needs to reach the designated goal, with a maximum climbing capability of $0.3$ meters as a constraint. We set the maximum simulation time to $120$s and conducted $10$ repeated trials for each scenario. A trial is considered successful if the robot achieves the goal within $120$s. Otherwise, the trial is considered a failure.
Evaluation Metrics. The evaluation metrics include success rate (SR), overall time (OT), overall time under successful trials (OTS), planning time (PT), execution time (ET), and trajectory length under successful trials (TLS) to comprehensively assess the performance of different methods. SR refers to the portion of successful trials across 10 trials. OT denotes the average simulation time across all trials. If a trial fails, OT is recorded as the maximum simulation time, which is $120$s. OTS represents the average task completion time for successful trials. PT and ET indicate the average time spent on task planning and skill execution under successful completion, respectively. TLS refers to the average trajectory length of successful trials. SR, OT, OTS, and TLS indicate the overall effectiveness and efficiency of methods, while PT and ET reflect the computational efficiency and quality of task planning.
@@
Figure 6: Success cases of AINav in maze environments. The green markers above the figures indicate the robot interacts with objects, while the red markers represent the robot performing replanning based on new environmental observations.
Figure 7: Success cases of AINav in Matterport3d, field ruin, and urban ruin.
V-B Baseline Methods
We compare the performance of AINav to the following baselines. Art-planner [wellhausen2023artplanner] is a sampling-based approach with learned motion cost for quadruped robots, chosen as a representative approach to solve traditional navigation tasks. Hierarchical RL trains a high-level policy to function as the task planning module, which decides the low-level skill to use and the corresponding parameters to reach the goal. We also select two LLM-based method for comparison. The first is RoboTool [xu2023creative], which employs multiple LLMs to perform different roles and invoke low-level skills to utilize tools, thereby solving complex planning and reasoning challenges. Since this method uses an open-loop formulation without reaction to new objects, we provide the complete environment description as its input and execute the skills suggested by its output. The second method is Tree-planner [hu2023tree], which employs a tree-structured reasoning approach to solve closed-loop TAMP problems. Because its plan sampling module requires global information, we include the initial object information as input. We adjust the prompts used in these two LLM-based methods for better adaptation to the interactive navigation tasks. For a fair comparison, all methods use the same set of low-level skills as AINav.
V-C Results and Analyses
Simulation results are displayed in Table˜I. We can observe that AINav exhibits higher success rates and less overall time compared to other baselines in diverse scenarios. The Art-planner achieves high task success rates in low-difficulty scenarios of the maze environment with low execution time, showcasing its effective planning capabilities in traditional navigation tasks. However, in other scenarios that require interaction, Art-planner is limited to passively adapting to the current environment without the capability to actively alter the workspace, thereby failing to accomplish more complex interactive navigation tasks. Hierarchical RL can achieve limited success in low-difficulty scenarios, while in more intricate tasks that require interaction to solve, hierarchical RL struggles to identify the precise sequence of skills necessary for task completion in the expansive search space.
@@ V-C Results and Analyses
Leveraging the scene understanding and reasoning capabilities of pre-trained LLM, RoboTool and tree-planner achieve a higher success rate in interactive navigation tasks compared to Art-planner. However, RoboTool is constrained by its open-loop planning scheme, which prevents it from adapting to new environmental information, resulting in lower success rates in high-difficulty tasks and disaster scenes. While tree-planner leverages a closed-loop design to adjust its decisions based on new observations, its performance is highly dependent on the quality of the sampled plans, resulting in instability in planning results. Specifically, once the tree structure is generated, its inflexibility can hinder the planner from obtaining the optimal solution that does not exist in the original tree or changing to an updated better strategy in response to new observations. Additionally, both methods require omniscient environmental information to generate task decomposition plans, rendering them ineffective in unknown environments. In contrast, AINav not only enhances the quality of task planning outcomes through the design of a proposer and evaluator but also leverages adaptive replanning to flexibly adapt to new environmental information. This ensures efficient planning and real-time responsiveness in unfamiliar environments without global environmental information as prior, resulting in high success rates in various scenarios. Although node evaluation slightly increases planning time, the improved task planning quality and flexible replanning mechanism enable AINav to achieve higher success rates, shorter overall completion time, and trajectory length compared to all baseline approaches. As shown in Fig.˜5, AINav demonstrates the highest success rate and the lowest overall completion time among all approaches and across various scenarios, showcasing its superior effectiveness in complex navigation tasks.
@@
We visualize the performance of AINav across different tasks and scenarios in Fig.˜6 and Fig.˜7. In the box obstruction and Matterport scenario, the robot demonstrates the ability to comprehend the blocked environment, move boxes to create free space to reach the target position. In the box usage, stair building, and ruin scenarios, the robot actively utilizes movable obstacles to construct stairs, enabling access to higher areas, which are unreachable for a quadruped robot under normal circumstances. Notably, AINav’s advanced reasoning is highlighted in the stair-building and urban ruin tasks, where the robot must sequentially arrange boxes of varying heights in the correct order to climb to the prohibitively high platform. The successful completion of these tasks demonstrates the capacity of AINav for complex tool-use reasoning and execution. Furthermore, we can observe that the robot can respond to updated information, such as when the box fails to move or the robot identifies new objects and tasks. By leveraging an efficient closed-loop system, AINav enhances the robot’s capability to adapt to incremental environmental comprehension, thereby facilitating effective task completion.
TABLE II: Failure study of AINav in simulations
Primary Category Sub-Category Rate
Task Timeouts Scene understanding error $17.2\%$
Incorrect task decomposition $34.5\%$
Skill execution unfinished $24.1\%$
Irrecoverable execution Intra-skill failure $10.3\%$
Skill transition failure $13.8\%$
We further conduct a failure study to analyze the current limitations of AINav. We categorize all failures observed during simulations into two primary types: task timeouts and irrecoverable execution failures, where task timeouts occur when the robot does not complete the task within the allocated time limit, and irrecoverable execution failures involve a critical error during primitive skill execution, such as the robot falling over. Table˜II summarizes the distribution of these failure modes across simulations. In summary, this failure study reveals that the primary bottlenecks for AINav lie in the efficiency of task planning and the seamless integration of low-level skills, highlighting clear directions for future improvement. Furthermore, the assumption of simplified visual processing and reliance on predefined constraints to determine skill executability limit AINav’s generalization in realistic scenarios, which will be the focus for future work to enhance robustness and adaptability.
V-D Ablation Studies
We conduct ablation studies in maze environments to compare AINav with the following four variants: 1) AINav-Single that only generates a single plan with LLM in the task planning module; 2) AINav-Skill that directly generates skills without incorporating object-centric reasoning; 3) AINav-No that removes the replanning module; 4) AINav-F that only replans when a failure is encountered. We evaluate these methods in the middle-difficulty and high-difficulty scenarios of box obstruction, box usage, and stair building in the maze environment, which need interaction for task completion, to compare the SR and OT. The results of ablation studies are presented in Fig.˜8. AINav-Single exhibits low success rates, as the single plan generated by the LLM is often flawed or inefficient due to inherent uncertainties of LLM, leading to task failures. Benefiting from the tree structure that enhances the robustness and planning quality, AINav-Skill achieves a reasonable success rate. However, it lacks a holistic understanding of object utilization in complex scenarios, leading to inexecutable task decomposition plans. AINav-No performs poorly because it lacks a replanning phase, making it unable to adapt to new observations. Conversely, AINav-F improves task success rates through failure recovery but tends to miss new objects that could offer more optimal solutions, thus often resulting in excessive time spent and task timeouts. By integrating tree-structured reasoning with the adaptive replanning module, AINav enhances its task-planning capabilities and adaptability to incremental environmental information, achieving the highest success rates and improved time efficiency.